Grade 12th passMechanics

a bullet when fired to a target has its velocity decreased to 25% after penetrating 45cm into it. then the additional thickness required to stop the bullet is...

Angel
10 Years agoGrade 12th pass

1 Answer

Carmel Emmanuel
10 Years ago
Final velocity =u/4
(u/4)2=u2-2a(0.45)   (1)
Now velocity to be 0 after hiting let additional thickeness be c
0=u^2 -2a(0.45+x)    (2)
From eq(1) substitute the value of a
Finally you get x+0.45=0.48
X=0.03=3 cm